he weekly wages of 100 employees are given below. Find D4 and D8. Also find the wage


above which we have the richest 35% employees, correct to the nearest rupee.


Weekly Wages Below 50 50-100 100-150 150-200 200-250 250-300


No. of employees 10 70 80 100 150 90


Hint: Richest 35% employees earn above P65.


Ans.: D4= 170, D8 = 246.66, Richest 35% employees earn above Rs. 222.


1
Expert's answer
2021-12-13T08:44:24-0500

Class Below 50 50-100 100-150 150-200 200-250 250-300

Frequency 10 70 80 100 150 90

Cf 10 80 160 260 410 500

"D4=l+w(\frac{4n/10-cf}{f})"

"D4=150+50(\frac{200-160}{100})"

=170

"D8=l+w(\frac{8n/10-cf}{f})"

"D8=200+50(\frac{400-260}{150})"

=246.667

"P65=l+w(\frac{65n/100-cf}{f})"

"P65=200+50(\frac{325-260}{150})"

=221.667